# Osmotic forces modify lipid membrane fluctuations

### v1

@@
Abstract. In hydrodynamic descriptions of lipid bilayers, the membrane is often approximated as being impermeable to the surrounding, solute-containing fluid. However, biological and in vitro lipid membranes are influenced by their permeability and the resultant osmotic forces—whose effects remain poorly understood. Here, we study the dynamics of a fluctuating, planar lipid membrane that is ideally selective: fluid can pass through it, while the electrically-neutral solutes cannot. We find that the canonical membrane relaxation mode, in which internal membrane forces are balanced by fluid drag, no longer exists over all wavenumbers. Rather, this mode only exists when it is slower than solute diffusion—corresponding to a finite range of wavenumbers. The well-known equipartition result quantifying the size of membrane undulations due to thermal perturbations is consequently limited in its validity to the aforementioned range. Moreover, this range shrinks as the membrane surface tension is increased, and above a critical tension the membrane mode vanishes. Our findings are relevant when interpreting experimental measurements of membrane fluctuations, especially in vesicles at moderate to high tensions.
@@ Introduction
Significance
Lipid membranes are ubiquitous in biological and in vitro systems. Though they are often approximated as impermeable, water can slowly pass through the membrane. In this work, we study the consequences of lipid membranes being permeable to water, but not solutes (such as sugars). We find that when solutes are fast relative to the membrane—which occurs at low surface tensions and intermediate wavelengths—membrane undulations are indistinguishable from the impermeable prediction. However, at large and small wavelengths or high tensions, solutes are slow relative to the membrane and the impermeable result does not hold. Our findings are relevant to the many experimental investigations that apply impermeable results across all wavelengths, and all tensions, in their characterization.
The present study investigates how planar lipid membranes act on, and are influenced by, their surrounding environment. Flat membranes are found in various biological settings, including the endoplasmic reticulum [4], Golgi complex [5], and neuromuscular junction [6]. In addition, theories of impermeable bilayer fluctuations about a flat surface are often used to describe the thermal undulations of giant unilamellar vesicles (GUVs). We begin by investigating an impermeable membrane, where the dynamics of the solute are not considered. By starting with a well-understood system (see e.g. Refs. [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]), we are better equipped to study our primary system of interest: the semipermeable membrane. The temporal evolution of the normal modes is solved for within the linear theory, and the relative time scale of membrane and solute relaxation is found to dictate the nature of the coupled dynamics. When solutes relax more quickly, the dynamics are indistinguishable from the impermeable scenario. It is then straightforward to present an effective Langevin equation including thermal perturbations, and obtain the magnitude of the resultant height fluctuations. In contrast, when solutes are slow relative to the membrane, the impermeable dynamics are unphysical and the resultant behavior is outside the scope of a normal mode analysis within a linear description. As a result, the character of thermal undulations—long assumed to be that of an impermeable membrane when analyzing fluctuating GUVs—is no longer known. Experimental consequences of our findings are highlighted, and avenues for future work are discussed.
@@
The position $\bm{x}$ of the perturbed membrane is given by $\bm{x}(x,y,t)=x\mkern 1.0mu\bm{e}_{x}+y\mkern 1.0mu\bm{e}_{y}+h(x,y,t)\mkern 1.0mu\bm{e}_{z}$, where $x,y\in[0,\ell_{\text{c}}]$ and $\lvert h\rvert\ll\ell_{\text{c}}$ by construction. For planar geometries with a stationary base state, the in-plane and out-of-plane membrane equations are decoupled [26, 27]. To solve for the time evolution of the membrane height, we require only the $z$-component of the linear momentum balance of the membrane—often called the shape equation and given by [27]
@@
where the wavevector $\bm{q}=(q_{x},q_{y})$ has corresponding wavenumber $q:=(q_{x}^{\,2}+q_{y}^{\,2})^{1/2}$, $\omega_{\bm{q}}$ is the associated relaxation frequency, and normal modes are assumed independent. Fluid unknowns are also decomposed into planar Fourier modes, with the $z$-dependence to be solved for. The fluid pressure, for example, is expressed as
@@
Equation (4), with the temporal ansatz $h\sim e^{-\omega t}$ in Eq. (2), reveals the dynamics of the membrane are analogous to that of the spring–mass–damper from introductory mechanics. Here $\rho_{\mkern-1.0mu\text{eff}}:=\rho_{\mkern-1.0mu\text{m}}+\mkern 1.0mu4\mkern 1.0mu\rho_{\mkern-1.0mu\text{f}}/q$ approximates the combined inertia of membrane and fluid, $4\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq$ is the well-known hydrodynamic drag from the surrounding fluid, and $E:=\lambda_{\text{c}}\mkern 1.0muq^{2}+\tfrac{1}{2}\mkern 1.0muk_{\text{b}}\mkern 1.0muq^{4}$ captures the energetics of small membrane fluctuations: $\tfrac{1}{2}\,\ell_{\text{c}}^{\,2}\mkern 1.0muE\mkern 1.0mu\lvert\hat{h}_{\bm{q}}^{\,2}\rvert$ is the energy of mode $\bm{q}$ to quadratic order [34]. Since inertial effects are small, Eq. (4) has the two approximate solutions
@@
$\kappa\mkern 1.0mu\big(v^{\pm}_{\mkern-1.0muz}-h_{\mkern-1.0mu,\mkern 1.0mut}\big)\,=\,k_{\text{B}}\theta\mkern 1.0mu\left\llbracket\mkern 1.0muc\mkern 1.0mu\right\rrbracket\,-\,\left\llbracket\mkern 1.0mup\mkern 1.0mu\right\rrbracket~,$ (8)
where $k_{\text{B}}$ is Boltzmann’s constant, $\theta$ is the absolute temperature, and $\left\llbracket\mkern 1.0muc\mkern 1.0mu\right\rrbracket:=c^{+}-c^{-}$ is the jump in solute concentration at the membrane surface. In the language of irreversible thermodynamics [44, 45], Eq. (8) presents a linear relationship between the thermodynamic driving force $(k_{\text{B}}\theta\mkern 1.0mu\left\llbracket\mkern 1.0muc\mkern 1.0mu\right\rrbracket-\left\llbracket\mkern 1.0mup\mkern 1.0mu\right\rrbracket)$ and the corresponding thermodynamic flux $(v^{\pm}_{\mkern-1.0muz}-h_{\mkern-1.0mu,\mkern 1.0mut})$, with the latter a measure of fluid flow through the membrane. The phenomenological parameter $\kappa$, understood as a transport coefficient [3], captures the resistance to flow: for a given driving force, a larger value of $\kappa$ corresponds to a smaller water flux through the membrane. Assuming the usual Stokes flow scaling of $\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq^{2}v\sim q\mkern 1.0mup$, in the limit of $\kappa\rightarrow\infty$ or $\mathcal{P}\rightarrow 0$ we recover the impermeable coupling condition $v^{\pm}_{\mkern-1.0muz}=h_{\mkern-1.0mu,\mkern 1.0mut}$.
@@
with $\tilde{\omega}_{\mkern-1.0mu\text{m}}<\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$ when $q\in(q_{0}^{-},q_{0}^{+})$, and $\tilde{\omega}_{\mkern-1.0mu\text{m}}>\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$ otherwise. Since the base surface tension $\lambda_{\text{c}}$ is not a material parameter, but rather varies—and can be tuned—across systems, $q_{0}^{\pm}$ is plotted at different tensions $\lambda_{\text{c}}$ in Fig. 2(b) to reveal a dome-like envelope. There exists a critical base tension
@@
at the top of the dome, above which no real $q_{0}$ exists. When the base membrane tension $\lambda_{\text{c}}$ is larger than $\lambda_{\text{c}}^{\mkern-1.0mu\ast}$, $\tilde{\omega}_{\mkern-1.0mu\text{m}}>\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$ over all wavenumbers. As we will see, the relative values of $\tilde{\omega}_{\mkern-1.0mu\text{m}}$ and $\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$—and thus whether a mode is “inside” (i.e. below) or “outside” (i.e. above) the dome in Fig. 2(b)—govern different regimes of membrane behavior.
@@
The inertial branch. We begin by assuming concentration modes evolve with the impermeable inertial frequency $\omega^{\text{imp}}_{\mkern-2.0mu\rho}$. The diffusion equation for mode $\bm{q}$ then simplifies to
$\dfrac{\text{d}^{2}\mkern 1.0mu{\hat{c}^{\mkern 1.0mu\pm}_{\bm{q}}}}{\text{d}\mkern 1.0mu{z}^{2}}\,=\,-q^{2}\mkern 1.0mu\tilde{\eta}^{2}\mkern 1.0mu\hat{c}^{\mkern 1.0mu\pm}_{\bm{q}}~,\qquad\text{where}\qquad\tilde{\eta}\,:=\mkern 1.0mu\bigg(\dfrac{\omega^{\text{imp}}_{\mkern-2.0mu\rho}}{\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}}\,-\,1\bigg)^{\!\!1\mkern-1.0mu/\mkern-1.0mu2}$ (13)
@@
We first recognize that water permeation alters the hydrodynamic drag on the membrane by the surrounding fluid. Equations (7) and (8) together yield $\left\llbracket\mkern 1.0mup\mkern 1.0mu\right\rrbracket=k_{\text{B}}\theta\mkern 1.0mu\left\llbracket\mkern 1.0muc\mkern 1.0mu\right\rrbracket-(\kappa\mkern 1.0muD/c_{0})\mkern 1.0muc^{\pm}_{\mkern-1.0mu,\mkern 1.0muz}\mkern 1.0mu$, which—along with Eq. (1)—shows standing waves in the concentration add a hydrodynamic drag force 90 out of phase with the membrane velocity. We show in the SM [25, §3.3.2] that an approximate dispersion relation for the semipermeable membrane is given by [cf. Eq. (4)]
@@
In Eq. (14), $\mathcal{S}:=k_{\text{B}}\theta\mkern 1.0muc_{0}/(D\mkern 1.0mu\kappa\mkern 1.0muq)\ll 1$ as a dimensionless parameter that arises when Eqs. (7) and (8) are combined. The solution of Eq. (14) closest to $\omega^{\text{imp}}_{\mkern-2.0mu\rho}$, to lowest order in the product $\mathcal{P}\mkern 1.0mu\mathcal{S}$ of small parameters, is given by
@@
Theoretical implications. As mentioned previously, solutes are present in the fluid surrounding nearly all artificial and biological membranes. Our findings demonstrate such systems can only be approximated as impermeable for wavenumbers “inside the dome,” where $q\in(q_{0}^{-},q_{0}^{+})$ and there is a slow frequency $\omega\approx E/(4\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq)$. In contrast, only a quickly decaying inertial mode exists for wavenumbers $q\notin(q_{0}^{-},q_{0}^{+})$ “outside the dome.” For this case, ($i$) the membrane response is independent of $k_{\text{b}}$ and $\lambda_{\text{c}}$ at linear order, and ($i\mkern-1.0mui$) thermal perturbations (discussed subsequently) drive membrane undulations to continuously grow until saturated by nonlinear forces—both of which are surprising results. Such predictions motivate us to question whether our normal mode ansatz is indeed appropriate outside the dome. If instead the solute concentration was decomposed as $c\sim\check{c}(z,t)\mkern 1.0mue^{i\mkern 1.0mu(q_{x}x\,+\,q_{y}y)}$ [cf. Eq. (12)], then the diffusion equation requires $\partial\check{c}/\partial t=D(\partial^{2}\check{c}/\partial z^{2}-q^{2}\check{c})$. Given an infinite domain $z\in(-\infty,\infty)$, a self-similar solution is expected for $\check{c}(z,t)$—which is incompatible with our ansatz. Further analysis, likely involving nonlinear simulations of the membrane and surrounding fluid [46, 47, 48, 49, 50, 51, 52], is required to understand the dynamics of the coupled system. In any case, however, we emphasize that the standard impermeable result is not applicable for wavenumbers outside the dome.
@@
In Eq. (20), $\tilde{\xi}_{\bm{q}}(t)$ is a complex Gaussian thermal noise satisfying the fluctuation–dissipation theorem, with two-time correlations proportional to $\delta(t-t^{\prime})\,k_{\text{B}}\theta/(4\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq\,\ell_{\text{c}}^{\,2})$ [18]. Equation (20) can be solved exactly to yield [60]
$\langle\lvert\mkern 1.0mu\tilde{h}_{\bm{q}}\rvert^{2}\rangle\,=\,\dfrac{k_{\text{B}}\theta}{E\mkern 1.0mu\ell_{\text{c}}^{\,2}}~,$ (21)
@@
Importantly, $k_{\text{b}}$ is not known a priori, and so Eq. (24) can be approximated as $\ell_{\text{c}}\mkern 1.0mu\langle\lvert\mkern 1.0mu\tilde{h}_{q_{x}}\mkern-1.0mu\rvert^{2}\rangle\approx k_{\text{B}}\theta/(8\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq_{x}{}^{\!\!2}D)$—valid when $k_{\text{b}}\mkern 1.0muq_{x}\ll 8\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muD$—in a first analysis of the data. Equation (24) is plotted as the dashed black line in Fig. 3(c); all of the data points below it are outside the dome and should be excluded when determining $k_{\text{b}}$ and $\lambda_{\text{c}}$.
@@
The vanishing of the membrane mode at low $q$ and large $\lambda_{\text{c}}$ opens several lines of inquiry. We are particularly interested to see if membranes in spherical or cylindrical geometries respond in the same fashion; the former is needed for a careful analysis of GUV undulations [56, 57, 58] while the latter could be relevant to neuronal systems [61, 30, 63, 62, 64]. In all scenarios where the membrane branch vanishes, the behavior of the coupled membrane–fluid–solute system can be interrogated with nonlinear simulations [46, 47, 48, 49, 50, 51, 52]. Additional levels of complexity can be introduced by considering membranes that are not ideally selective [3] or charged solutes—with the latter requiring extensive theoretical developments [73, 74, 75, 76, 77].
@@
Acknowledgments. We express our most sincere gratitude to Prof. Kranthi Mandadapu for his perspective on lipid membrane behavior. We are also thankful to have discussed Ref. [3] with Dr. Ahmad Alkadri and Kranthi, as well as Refs. [59, 55] with Dr. Kisung Lee, Dr. Myoenggon Park, and Prof. Steve Granick. It is a pleasure to thank Dr. Joël Tchoufag for many stimulating discussions, Profs. Brian Camley and Tine Curk regarding their thoughts on membrane fluctuations, and Prof. Sho Takatori for introducing us to the experimental analysis of GUV fluctuations.

### v2

@@
Abstract. In hydrodynamic descriptions of lipid bilayers, the membrane is often approximated as being impermeable to the surrounding, solute-containing fluid. However, biological and in vitro lipid membranes are influenced by their permeability and the resultant osmotic forces—whose effects remain poorly understood. Here, we study the dynamics of a fluctuating, planar lipid membrane that is ideally selective: fluid can pass through it, while solutes cannot. We find that the canonical membrane relaxation mode, in which internal membrane forces are balanced by fluid drag, no longer exists over all wavenumbers. Rather, this mode only exists when it is slower than solute diffusion—corresponding to a finite range of wavenumbers. The well-known equipartition result quantifying the size of membrane undulations due to thermal perturbations is consequently limited in its validity to the aforementioned range. Moreover, this range shrinks as the membrane surface tension is increased, and above a critical tension the membrane mode vanishes. Our findings are relevant when interpreting experimental measurements of membrane fluctuations, especially in vesicles at moderate to high tensions.
@@ Introduction
The present study investigates how planar lipid membranes act on, and are influenced by, their surrounding environment. Flat membranes are found in various biological settings, including the endoplasmic reticulum [4], Golgi complex [5], and neuromuscular junction [6]. In addition, theories of impermeable bilayer fluctuations about a flat surface are often used to describe the thermal undulations of giant unilamellar vesicles (GUVs). We begin by investigating an impermeable membrane, where the dynamics of the solute are not considered. By starting with a well-understood system (see e.g. Refs. [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]), we are better equipped to study our primary system of interest: the semipermeable membrane. The temporal evolution of the normal modes is solved for within the linear theory, and the relative time scale of membrane and solute relaxation is found to dictate the nature of the coupled dynamics. When solutes relax more quickly, the dynamics are indistinguishable from the impermeable scenario. It is then straightforward to present an effective Langevin equation including thermal perturbations, and obtain the magnitude of the resultant height fluctuations. In contrast, when the membrane relaxes more quickly than the surrounding solutes, the resultant dynamics are qualitatively different from their impermeable counterpart. As a result, the character of thermal undulations—long assumed to be that of an impermeable membrane when analyzing fluctuating GUVs—is no longer known. Experimental consequences of our findings are highlighted, and avenues for future work are discussed.
@@
The position $\bm{x}$ of the perturbed membrane is given by $\bm{x}(x,y,t)=x\mkern 1.0mu\bm{e}_{x}+y\mkern 1.0mu\bm{e}_{y}+h(x,y,t)\mkern 1.0mu\bm{e}_{z}$, where $x,y\in[0,\ell_{\text{c}}]$ and $\lvert h\rvert\ll\ell_{\text{c}}$ by construction. For planar geometries with a stationary base state, the in-plane and out-of-plane membrane equations are decoupled [26, 27]. To solve for the time evolution of the membrane height in the absence of thermal fluctuations, we require only the $z$-component of the linear momentum balance of the membrane—often called the shape equation and given by [27]
@@
where the wavevector $\bm{q}=(q_{x},q_{y})$ has corresponding wavenumber $q:=(q_{x}^{\,2}+q_{y}^{\,2})^{1/2}$ $\sim 10^{-4}\text{--}10^{-1}\text{ nm}^{-1}\!\!$,  $\omega_{\bm{q}}$ is the associated relaxation frequency, and normal modes are assumed independent. Fluid unknowns are similarly decomposed into planar Fourier modes, with the $z$-dependence to be solved for. The fluid pressure, for example, is expressed as
@@
[…]
@@
$\kappa\mkern 1.0mu\big(v^{\pm}_{\mkern-1.0muz}-h_{\mkern-1.0mu,\mkern 1.0mut}\big)\,=\,k_{\text{B}}\vartheta\mkern 1.0mu\left\llbracket\mkern 1.0muc\mkern 1.0mu\right\rrbracket\,-\,\left\llbracket\mkern 1.0mup\mkern 1.0mu\right\rrbracket~,$ (8)
where $k_{\text{B}}$ is Boltzmann’s constant, $\vartheta$ is the absolute temperature, and $\left\llbracket\mkern 1.0muc\mkern 1.0mu\right\rrbracket:=c^{+}-c^{-}$ is the jump in solute concentration at the membrane surface. In the language of irreversible thermodynamics [44, 45], Eq. (8) presents a linear relationship between the thermodynamic driving force $(k_{\text{B}}\vartheta\mkern 1.0mu\left\llbracket\mkern 1.0muc\mkern 1.0mu\right\rrbracket-\left\llbracket\mkern 1.0mup\mkern 1.0mu\right\rrbracket)$ and the corresponding thermodynamic flux $(v^{\pm}_{\mkern-1.0muz}-h_{\mkern-1.0mu,\mkern 1.0mut})$, with the latter a measure of fluid flow through the membrane. The phenomenological parameter $\kappa$, understood as a transport coefficient [3], captures the resistance to flow: for a given driving force, a larger value of $\kappa$ corresponds to a smaller water flux through the membrane. Equation (8) is equivalently expressed as $v^{\pm}_{\mkern-1.0muz}-h_{\mkern-1.0mu,\mkern 1.0mut}=(k_{\text{B}}\vartheta\mkern 1.0mu\left\llbracket\mkern 1.0muc\mkern 1.0mu\right\rrbracket-\left\llbracket\mkern 1.0mup\mkern 1.0mu\right\rrbracket)/\kappa$; in the limit of $\kappa\rightarrow\infty$ or $\mathcal{P}\rightarrow 0$ we find $v^{\pm}_{\mkern-1.0muz}\rightarrow h_{\mkern-1.0mu,\mkern 1.0mut}$ and recover the impermeable scenario.
@@
with $\tilde{\omega}_{\mkern-1.0mu\text{m}}<\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$ when $q\in(q_{0}^{-},q_{0}^{+})$, and $\tilde{\omega}_{\mkern-1.0mu\text{m}}>\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$ otherwise. Since the base surface tension $\lambda_{\text{c}}\sim 10^{-4}\text{--}10^{-1}$ pN/nm is not a material parameter, but rather varies—and can be tuned—across systems, $q_{0}^{\pm}$ is plotted at different tensions $\lambda_{\text{c}}$ in Fig. 2(b) to reveal a dome-like envelope. There exists a critical base tension
@@
at the top of the dome, above which no real $q_{0}$ exists. We also note that the bilayer tears at the lysis tension $\lambda_{\ell}\sim 1\text{--}10$ pN/nm [1]. Thus, when the base membrane tension $\lambda_{\text{c}}$ lies between $\lambda_{\text{c}}^{\mkern-1.0mu\ast}$ and $\lambda_{\ell}$, we have $\tilde{\omega}_{\mkern-1.0mu\text{m}}>\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$ over all wavenumbers. As we will see, the relative values of $\tilde{\omega}_{\mkern-1.0mu\text{m}}$ and $\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}$—and thus whether a mode is “inside” (i.e. below) or “outside” (i.e. above) the dome in Fig. 2(b)—govern different regimes of membrane behavior.
@@
The inertial branch. We begin by assuming concentration modes evolve with the impermeable inertial frequency $\omega^{\text{imp}}_{\mkern-2.0mu\rho}$, such that for a single mode, $c^{\pm}_{,t}=-\omega^{\text{imp}}_{\mkern-2.0mu\rho}\mkern 1.0muc^{\pm}$ (see Eq. (12)). As a consequence, the diffusion equation $c^{\pm}_{,t}=D\mkern 1.0muc^{\pm}_{,jj}$ for the mode $\bm{q}$ simplifies to $\omega^{\text{imp}}_{\mkern-2.0mu\rho}\mkern 1.0mu\hat{c}^{\mkern 1.0mu\pm}_{\bm{q}}=D(-q^{2}\mkern 1.0mu\hat{c}^{\mkern 1.0mu\pm}_{\bm{q}}+\text{d}^{2}\hat{c}^{\mkern 1.0mu\pm}_{\bm{q}}{}/\text{d}z^{2})$—which after some rearrangement yields
$\dfrac{\text{d}^{2}\mkern 1.0mu{\hat{c}^{\mkern 1.0mu\pm}_{\bm{q}}}}{\text{d}\mkern 1.0mu{z}^{2}}\,=\,-q^{2}\mkern 1.0mu\tilde{\eta}^{2}\mkern 1.0mu\hat{c}^{\mkern 1.0mu\pm}_{\bm{q}}~.\qquad\text{Here}\qquad\tilde{\eta}\,:=\mkern 1.0mu\bigg(\dfrac{\omega^{\text{imp}}_{\mkern-2.0mu\rho}}{\tilde{\omega}_{\mkern-2.0mu\scriptscriptstyle D}}\,-\,1\bigg)^{\!\!1\mkern-1.0mu/\mkern-1.0mu2}$ (13)
@@
We first recognize that water permeation alters the hydrodynamic drag on the membrane by the surrounding fluid. Equations (7) and (8) together yield $\left\llbracket\mkern 1.0mup\mkern 1.0mu\right\rrbracket=k_{\text{B}}\vartheta\mkern 1.0mu\left\llbracket\mkern 1.0muc\mkern 1.0mu\right\rrbracket-(\kappa\mkern 1.0muD/c_{0})\mkern 1.0muc^{\pm}_{\mkern-1.0mu,\mkern 1.0muz}\mkern 1.0mu$, which—along with Eq. (1)—shows standing waves in the concentration add a hydrodynamic drag force 90 out of phase with the membrane velocity. We show in the SM [25, §3.3.2] that an approximate dispersion relation for the semipermeable membrane is given by [cf. Eq. (4)]
@@
In Eq. (14), $\mathcal{S}:=k_{\text{B}}\vartheta\mkern 1.0muc_{0}/(D\mkern 1.0mu\kappa\mkern 1.0muq)$ $\sim 10^{-8}\text{--}10^{-3}\ll 1$ is a dimensionless parameter that arises when Eqs. (7) and (8) are combined. The solution of Eq. (14) closest to $\omega^{\text{imp}}_{\mkern-2.0mu\rho}$, to lowest order in the product $\mathcal{P}\mkern 1.0mu\mathcal{S}$ of small parameters, is given by
@@
Theoretical implications. As mentioned previously, solutes are present in the fluid surrounding nearly all artificial and biological membranes. Our findings demonstrate such systems can only be approximated as impermeable for wavenumbers “inside the dome,” where $q\in(q_{0}^{-},q_{0}^{+})$ and there is a slow frequency $\omega\approx E/(4\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq)$. In contrast, only a quickly decaying inertial mode exists for wavenumbers $q\notin(q_{0}^{-},q_{0}^{+})$ “outside the dome.” For this case, ($i$) the membrane response is independent of $k_{\text{b}}$ and $\lambda_{\text{c}}$ at linear order, and ($i\mkern-1.0mui$) thermal perturbations (discussed subsequently) drive membrane undulations to continuously grow until saturated by nonlinear forces—both of which are surprising results.
The predictions described above motivate us to question whether our normal mode ansatz is indeed appopriate outside the dome. Here, we examine this issue from a statistical mechanical perspective. By starting with a microscopic Hamiltonian of the lipid, water, and solute molecules, one can integrate over fast degrees of freedom to obtain a free energy functional that depends only on the concentration field and membrane height [46]. When solutes are fast relative to the membrane, concentration degrees of freedom can also be integrated over to yield a free energy which depends only on the membrane height. We thus recover a description similar to that of the impermeable scenario. If instead solutes are slow relative to the membrane, the concentration field is not equilibrated for a given, instantaneous bilayer shape. Upon integrating over the slow solute degrees of freedom, the resultant membrane equation inerits this slowness via a memory kernel—for which the time evolution of the membrane depends on its history, and there is no associated free energy functional or equipartition result [47]. While such a result is incompatible with our normal mode ansatz, it justifies the membrane behaving qualitatively differently inside and outside the dome. Further analysis, likely involving nonlinear simulations of the membrane and surrounding fluid [48, 49, 50, 51, 52, 53, 54], is required to understand the dynamics of the coupled system. In any case, however, we emphasize that the standard impermeable result is not applicable for wavenumbers outside the dome.
@@
In Eq. (20), $\tilde{\xi}_{\bm{q}}(t)$ is a complex Gaussian thermal noise satisfying the fluctuation–dissipation theorem, with two-time correlations proportional to $\delta(t-t^{\prime})\,k_{\text{B}}\vartheta/(4\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq\,\ell_{\text{c}}^{\,2})$ [18]. Equation (20) can be solved exactly to yield [62]
$\langle\lvert\mkern 1.0mu\tilde{h}_{\bm{q}}\rvert^{2}\rangle\,=\,\dfrac{k_{\text{B}}\vartheta}{E\mkern 1.0mu\ell_{\text{c}}^{\,2}}~,$ (21)
@@
Importantly, $k_{\text{b}}$ is not known a priori, and so Eq. (24) can be approximated as $\ell_{\text{c}}\mkern 1.0mu\langle\lvert\mkern 1.0mu\tilde{h}_{q_{x}}\mkern-1.0mu\rvert^{2}\rangle\approx k_{\text{B}}\vartheta/(8\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muq_{x}{}^{\!\!2}D)$—valid when $k_{\text{b}}\mkern 1.0muq_{x}\ll 8\mkern 1.0mu\mu_{\mkern-1.0mu\text{f}}\mkern 1.0muD$—in a first analysis of the data. Equation (24) is plotted as the dashed black line in Fig. 3(c); all of the data points below it are outside the dome and should be excluded when determining $k_{\text{b}}$ and $\lambda_{\text{c}}$.
@@
The vanishing of the membrane mode at low $q$ and large $\lambda_{\text{c}}$ opens several lines of inquiry. We are particularly interested to see if membranes in spherical or cylindrical geometries respond in the same fashion; the former is needed for a careful analysis of GUV undulations [58, 59, 60] while the latter could be relevant to neuronal systems [63, 30, 65, 64, 66]. In all scenarios where the membrane branch vanishes, the behavior of the coupled membrane–fluid–solute system can be interrogated with nonlinear simulations [48, 49, 50, 51, 52, 53, 54]. Additional levels of complexity can be introduced by considering membranes that are not ideally selective [3], have finite thickness [75, 76, 77, 78], or are surrounded by charged solutes—with the latter requiring extensive theoretical developments [79, 76, 77, 78, 80].
@@
Acknowledgments. We express our most sincere gratitude to Prof. Kranthi Mandadapu for his perspective on lipid membrane behavior. We are also thankful to have discussed Ref. [3] with Dr. Ahmad Alkadri and Kranthi, as well as Refs. [61, 57] with Dr. Kisung Lee, Dr. Myoenggon Park, and Prof. Steve Granick. It is a pleasure to thank Dr. Joël Tchoufag for many stimulating discussions, Profs. Brian Camley and Tine Curk regarding their thoughts on membrane fluctuations, and Prof. Sho Takatori for introducing us to the experimental analysis of GUV fluctuations. This work was partially supported by the Welch Foundation via Grant No. F-2208-20240404.
Declaration of interests. There are no conflicts to declare.